# Methods and stability tests associated with the sterile neutrino search using improved high-energy $\nu_{\mu}$ event reconstruction in IceCube

### v1

@@ VI Conclusion
This paper has presented detailed information related to the first search for a $\nu_{\mu}$ disappearance signature due to a 3+1 model in IceCube that separately fits 11 years of data divided into starting and through-going samples. This analysis makes use of neutrino events with reconstructed energy in the 0.5 to 100 TeV region and upward going zenith angle ($\cos(\theta_{z}^{\mathrm{reco}})<0$). The result is reported in Ref. [7].
@@
We presented an extensive discussion of the expected signature, focusing on the bin-by-bin predictions from the best fit, which was at the parameters $\sin^{2}(2\theta_{24})=0.16$ and $\Delta m^{2}_{41}=3.5$ eV2. The p-value representing the goodness of fit, found by trials, is 12%. Consistency with the null hypothesis of no sterile neutrino oscillations has a probability of 3.1%. In this article, we have also presented a Bayesian analysis. While credible regions have a different meaning than confidence limits, the overlap of the best-fit parameter space indicates the consistency of the result even under various approaches.

### v2

@@ VI Conclusion
This paper has presented detailed information related to the first search for a $\nu_{\mu}$ disappearance signature due to a 3+1 model in IceCube that separately fits 10.7 years of data divided into starting and through-going samples. This analysis makes use of neutrino events with reconstructed energy in the 0.5 to 100 TeV region and upward going zenith angle ($\cos(\theta_{z}^{\mathrm{reco}})<0$). The result is reported in Ref. [7].
@@
We presented an extensive discussion of the expected signature, focusing on the bin-by-bin predictions from the best fit, which was at the parameters $\sin^{2}(2\theta_{24})=0.16$ and $\Delta m^{2}_{41}=3.5$ eV2. The p-value representing the goodness of fit, found by trials, is 12%. Consistency with the null hypothesis of no sterile neutrino oscillations has a probability of 3.1%. In this article, we have also presented a Bayesian analysis. While the Bayesian model comparison has a different meaning than the frequentist result, the overlap of the best-fit parameter space indicates the consistency of the result even under various approaches.